# The lower $p$-series of analytic pro-$p$ groups and Hausdorff dimension

### v2

@@ 2. Stratifications of the lower $p$-series
We proceed to establish three auxiliary claims. As on previous occasions we embed $L$ into $\mathbb{Q}_{p}\otimes_{\mathbb{Z}_{p}}L$ to interpret terms involving negative powers of $p$ and lying outside $L$.
Claim 1. There are elements $w_{i,k}\in N$, for $i\in\mathbb{N}_{0}$ and $\tilde{e}<k\leq e$, and $n_{i}\in\mathbb{N}$ with $n_{i}\leq\lfloor i\eta\rfloor$, for $i\in\mathbb{N}_{0}$, such that $L_{i}$, $i\in\mathbb{N}_{0}$, is equivalent to the series
$\bigoplus_{k=1}^{\tilde{e}}p^{\lfloor i\xi_{k}\rfloor}\mathbb{Z}_{p}x_{k}\;\oplus\;\bigoplus_{k=\tilde{e}+1}^{e}\mathbb{Z}_{p}\big{(}p^{\lfloor i\xi_{k}\rfloor}x_{k}+w_{i,k}\big{)}\;\oplus\;p^{n_{i}}N,\quad i\in\mathbb{N}_{0}.$ (2.5)
@@
Moreover, the $G$-invariant $\mathbb{Z}_{p}$-lattice $\tilde{M}=\bigoplus_{k=1}^{\tilde{e}}\mathbb{Z}_{p}x_{k}$ embeds isomorphically as $(\tilde{M}+N)/N$ into $L/N$. By Proposition 2.4, applied to $L/N$, there exists $c_{2}\in\mathbb{N}$ such that for all $i\in\mathbb{N}_{0}$,
$p^{c_{2}}\bigoplus_{k=1}^{\tilde{e}}p^{\lfloor i\xi_{k}\rfloor}\mathbb{Z}_{p}x_{k}\subseteq\lambda_{i}(\tilde{M})\subseteq L_{i}.$
@@
$p^{c_{3}}p^{\lfloor i\xi_{k}\rfloor}x_{k}\in L_{i}\qquad\text{for $1\leq k\leq\tilde{e}$.}$
@@
allows us to select $w_{i,\tilde{e}+1},\ldots,w_{i,e}\in N$ such that
$p^{c_{3}}\big{(}p^{\lfloor i\xi_{k}\rfloor}x_{k}+w_{i,k}\big{)}\in L_{i}\qquad\text{for $\tilde{e}<k\leq e$.}$
For $i\in\mathbb{N}_{0}$ with $i<j$ we set $w_{i,\tilde{e}+1}=\ldots=w_{i,e}=0$; small values of $i$ play an insignificant role and this simple choice has no relevance beyond the fact that the series in (2.5) is properly defined. To streamline the notation, we also set
$w_{i,1}=\ldots=w_{i,\tilde{e}}=0\qquad\text{for all $i\in\mathbb{N}_{0}$.}$ (2.6)
@@
We have established that $L_{i}$, $i\in\mathbb{N}_{0}$, is equivalent to the filtration series
@@
By Proposition 2.2, the filtration series $L_{i}\cap N$, $i\in\mathbb{N}_{0}$, of $N$ is equivalent to a filtration series of the form $p^{n_{i}}N$, $i\in\mathbb{N}_{0}$, and our choice of $\eta$, in connection with the lower $p$-series of $N$, yields $c_{4}\in\mathbb{N}$ such that $p^{\lfloor i\eta\rfloor+c_{4}}N\subseteq\lambda_{i}(N)\subseteq L_{i}$ for $i\in\mathbb{N}_{0}$. This allows us to arrange that $n_{i}\in\mathbb{N}$ satisfies $n_{i}\leq\lfloor i\eta\rfloor$ for all $i\in\mathbb{N}_{0}$. Recalling the notational ‘nullnummer’ (2.6), we deduce that $L_{i}$, $i\in\mathbb{N}_{0}$, is equivalent to the filtration series in (2.5), as claimed. $\diamond$
Claim 2. There exists $b\in\mathbb{N}$ such that for all $i\in\mathbb{N}_{0}$,
$L_{i}\subseteq M\oplus p^{\min\{\lfloor i\vartheta\rfloor,\lfloor i\eta\rfloor\}-b}N.$ (2.7)
Clearly, the $\mathbb{Z}_{p}$-lattice $L$ decomposes as a direct sum
$L=\tilde{M}\oplus K\oplus N$
of its sublattices $\tilde{M}$, $K=\bigoplus_{k=\tilde{e}+1}^{e}\mathbb{Z}_{p}x_{k}$ and $N$. Furthermore, the lattices $\tilde{M}$ and $N$ are $G$-invariant, whereas $K$ is not.
Recall that $\mathfrak{a}_{G}\trianglelefteq\mathbb{Z}_{p}G$ denotes the $p$-augmentation ideal and yields the description $L_{i}=L.\mathfrak{a}_{G}^{\,i}$, $i\in\mathbb{N}_{0}$, of the lower $p$-series of $L$, and similarly for $N$. According to our set-up, there exists $c_{1}\in\mathbb{N}$ such that for all $i\in\mathbb{N}_{0}$,
@@
$p^{c_{1}}\big{(}K.\mathfrak{a}_{G}^{\,i}\big{)}\subseteq p^{c_{1}}L_{i}\subseteq\tilde{M}\oplus p^{\lfloor i\vartheta\rfloor}K\oplus N\qquad\text{and}\qquad p^{c_{1}}\big{(}N.\mathfrak{a}_{G}^{\,i}\big{)}\subseteq p^{\lfloor i\eta\rfloor}N.$
By decomposing elements into their $\tilde{M}$-, $K$- and $N$-coordinates and tracing their ‘trajectory’ as we descend along the lower $p$-series, we can make the inclusion recorded above more precise. We choose to work modulo $\tilde{M}$ and can thus focus on the $K$- and $N$-coordinates. Of significance are the $K$-coordinates, because they can give rise to non-trivial $K$- and $N$-coordinates at the next step. We introduce a parameter $j$, which indicates that we follow the $K$-coordinates for $j$ steps, then switch to $N$-coordinates and continue to follow the latter for the remaining $i-j$ steps. Concretely, we deduce that for all $i\in\mathbb{N}_{0}$,
$K.\mathfrak{a}_{G}^{\,i}$ $\subseteq\tilde{M}\oplus p^{\lfloor i\vartheta\rfloor-c_{1}}K\oplus\sum_{j=0}^{i-1}p^{\lfloor j\vartheta\rfloor-c_{1}}N.\mathfrak{a}_{G}^{\,i-1-j}$
$\subseteq\tilde{M}\oplus p^{\lfloor i\vartheta\rfloor-c_{1}}K\oplus\sum_{j=0}^{i-1}p^{\lfloor j\vartheta\rfloor+\lfloor(i-1-j)\eta\rfloor-2c_{1}}N$
$\subseteq\tilde{M}\oplus p^{\lfloor i\vartheta\rfloor-c_{1}}K\oplus p^{\min\{\lfloor(i-1)\vartheta\rfloor,\lfloor(i-1)\eta\rfloor\}-2c_{1}-2}N.$
This yields, for $i\in\mathbb{N}_{0}$,
$L_{i}=\tilde{M}.\mathfrak{a}_{G}^{\,i}+K.\mathfrak{a}_{G}^{\,i}+N.\mathfrak{a}_{G}^{\,i}\subseteq M+K.\mathfrak{a}_{G}^{\,i}+N.\mathfrak{a}_{G}^{\,i}\\ \subseteq M+K.\mathfrak{a}_{G}^{\,i}+p^{\lfloor i\eta\rfloor-c_{1}}N\subseteq M\oplus p^{\min\{\lfloor i\vartheta\rfloor,\lfloor i\eta\rfloor\}-2c_{1}-3}N.$
Thus (2.7) holds for $b=2c_{1}+3$. $\diamond$
Claim 3. Let $\mathcal{V}=\big{(}p^{-b}N\big{)}^{\hat{e}-\tilde{e}}$, with $b$ as in Claim 2. Then there exists $m\in\mathbb{N}_{0}$ such that for every choice of $\mathbf{v}=(v_{\tilde{e}+1},\ldots,v_{\hat{e}})\in\mathcal{V}$, we have
$p^{m}N\subseteq\tilde{M}+\sum_{k=\tilde{e}+1}^{\hat{e}}(x_{k}+v_{k}).\mathbb{Z}_{p}G.$ (2.8)
For each $\mathbf{v}=(v_{\tilde{e}+1},\ldots,v_{\hat{e}})\in\mathcal{V}$, we think of $\tilde{T}_{\mathbf{v}}=\sum_{k=\tilde{e}+1}^{\hat{e}}\mathbb{Z}_{p}(x_{k}+v_{k})$ as a deformation of $T$ inside $M\oplus p^{-b}N\subseteq\mathbb{Q}_{p}\otimes_{\mathbb{Z}_{p}}L$. Recall that, by our choice of $\tilde{e}$, the $\mathbb{Z}_{p}$-lattice $\tilde{M}\oplus p^{b}\tilde{T}_{\mathbf{v}}\subseteq L$ is not $G$-invariant, whereas $\tilde{M}\oplus p^{b}\tilde{T}_{\mathbf{v}}\oplus N=\tilde{M}\oplus p^{b}T\oplus N$ is $G$-invariant by Proposition 2.4. This implies that
$\big{(}\tilde{M}+\tilde{T}_{\mathbf{v}}.\mathbb{Z}_{p}G\big{)}\cap N\neq\{0\}.$
Clearly, the function $\mathcal{V}\to\mathbb{N}_{0}$, $\mathbf{v}\mapsto l_{\mathbf{v}}$ is locally constant and, since the domain $\mathcal{V}$ is compact, the maximum $m=\max\{l_{\mathbf{v}}\mid\mathbf{v}\in\mathcal{V}\}$ exists and has the property described in (2.8). $\diamond$
Now we just need to put the pieces together. Combining the assertions of Claim 1 and Claim 2, we see that there exists $c_{1}\in\mathbb{N}$ such that the lower $p$-series $L_{i}$, $i\in\mathbb{N}_{0}$, is $c_{1}$-equivalent to the filtration series
$\bigoplus_{k=1}^{\tilde{e}}p^{\lfloor i\xi_{k}\rfloor}\mathbb{Z}_{p}x_{k}\;\oplus\;\bigoplus_{k=\tilde{e}+1}^{e}\mathbb{Z}_{p}\big{(}p^{\lfloor i\xi_{k}\rfloor}x_{k}+w_{i,k}\big{)}\;\oplus\;p^{n_{i}}N,\quad i\in\mathbb{N}_{0},$
where
$w_{i,k}\in p^{\min\{\lfloor i\vartheta\rfloor,\lfloor i\eta\rfloor\}-b}N\quad\text{and}\quad\min\{\lfloor i\vartheta\rfloor,\lfloor i\eta\rfloor\}-b\leq n_{i}\leq\lfloor i\eta\rfloor$
for all relevant indices $i,k$.
First consider the case $\eta\leq\vartheta$. In this situation
@@
for all relevant indices $i,k$ and it follows that $L_{i}$, $i\in\mathbb{N}_{0}$, is $(b+c_{1})$-equivalent to the series
$\bigoplus_{k=1}^{\tilde{e}}p^{\lfloor i\xi_{k}\rfloor}\mathbb{Z}_{p}x_{k}\;\oplus\;\bigoplus_{k=\tilde{e}+1}^{e}\mathbb{Z}_{p}p^{\lfloor i\xi_{k}\rfloor}x_{k}\;\oplus\;p^{\lfloor i\eta\rfloor}N,\quad i\in\mathbb{N}_{0},$
@@
and, if $k\in\{1,\ldots,\tilde{e}\}$ is minimal subject to $\eta\leq\xi_{k}$, this yields a stratification with frame
@@
Using Claim 3, we deduce that for $i\in\mathbb{N}_{0}$,
$L_{i}\supseteq p^{c_{1}}\Big{(}p^{\lfloor i\vartheta\rfloor}\tilde{M}+\bigoplus_{k=\tilde{e}+1}^{\hat{e}}p^{\lfloor i\vartheta\rfloor}\mathbb{Z}_{p}(x_{k}+v_{i,k})\Big{)}.\mathbb{Z}_{p}G\\ \supseteq p^{c_{1}+\lfloor i\vartheta\rfloor}\Big{(}\tilde{M}+\sum_{k=\tilde{e}+1}^{\hat{e}}(x_{k}+v_{i,k}).\mathbb{Z}_{p}G\Big{)}\supseteq p^{c_{1}+m+\lfloor i\vartheta\rfloor}N$
and hence
$p^{n_{i}}N\supseteq p^{c_{1}}L_{i}\cap N\supseteq p^{2c_{1}+m+\lfloor i\vartheta\rfloor}N.$
It follows that $L_{i}$, $i\in\mathbb{N}_{0}$, is $(b+2c_{1}+m)$-equivalent to the series
$\bigoplus_{k=1}^{\tilde{e}}p^{\lfloor i\xi_{k}\rfloor}\mathbb{Z}_{p}x_{k}\;\oplus\;\bigoplus_{k=\tilde{e}+1}^{e}p^{\lfloor i\xi_{k}\rfloor}\mathbb{Z}_{p}x_{k}\;\oplus\;p^{\lfloor i\vartheta\rfloor}N,\quad i\in\mathbb{N}_{0}.$
@@
$(x_{1},x_{2},\ldots,x_{\tilde{e}},y_{1},y_{2},\ldots,y_{d-e},x_{\tilde{e}+1},x_{\tilde{e}+2},\ldots,x_{e})$
@@
$(\xi_{1},\xi_{2},\ldots,\xi_{\tilde{e}},\underbrace{\vartheta,\vartheta,\ldots,\vartheta}_{\text{$d-e$ entries}},\xi_{\tilde{e}+1},\xi_{\tilde{e}+2},\ldots,\xi_{e})\in\big{(}[\nicefrac{{1}}{{\max\{e,d-e\}}},1]\cap\mathbb{Q}\big{)}^{d}.\qed$
@@
Furthermore, for every $i\in\mathbb{N}$ with $i\geq j$, the powerfully embedded subgroup $P_{i}(G)\trianglelefteq_{\mathrm{o}}U$ gives, on the Lie lattice side, a powerfully embedded Lie sublattice $L_{i^{*}}\trianglelefteq_{\mathrm{o}}L$, where we write $i^{*}=i-j$ to ease notationally a recurrent constant shift by $j$ in the index; see [15, Thm. 1.4].

### v3

@@ 2. Stratifications of the lower $p$-series
As on previous occasions we embed $L$ into $\mathbb{Q}_{p}\otimes_{\mathbb{Z}_{p}}L$ to interpret terms involving negative powers of $p$ and lying outside $L$. The next step is to show that there exists $b_{0}\in\mathbb{N}$ such that
$L_{i}\subseteq M\oplus p^{\min\{\lfloor i\vartheta\rfloor,\lfloor i\eta\rfloor\}-b_{0}}N\qquad\text{for all $i\in\mathbb{N}_{0}$.}$ (2.5)
Indeed, the $\mathbb{Z}_{p}$-lattice $L$ clearly decomposes as a direct sum
$L=\tilde{M}\oplus K\oplus N$ (2.6)
of its sublattices $\tilde{M}$, $K=\bigoplus_{k=\widetilde{e}+1}^{e}\mathbb{Z}_{p}x_{k}$ and $N$. Furthermore, the lattices $\tilde{M}$ and $N$ are $G$-invariant, whereas $K$ is not.
Recall that $\mathfrak{a}_{G}\trianglelefteq\mathbb{Z}_{p}G$ denotes the $p$-augmentation ideal and yields the description $L_{i}=L.\mathfrak{a}_{G}^{\,i}$, $i\in\mathbb{N}_{0}$, of the lower $p$-series of $L$, and similarly for $N$. According to our set-up, there exists $b_{1}\in\mathbb{N}$ such that for all $i\in\mathbb{N}_{0}$,
$p^{b_{1}}\big{(}K.\mathfrak{a}_{G}^{\,i}\big{)}\subseteq p^{b_{1}}L_{i}\subseteq\tilde{M}\oplus p^{\lfloor i\vartheta\rfloor}K\oplus N\qquad\text{and}\qquad p^{b_{1}}\big{(}N.\mathfrak{a}_{G}^{\,i}\big{)}\subseteq p^{\lfloor i\eta\rfloor}N.$
By decomposing elements $y\in L$ according to (2.6) as $y=y_{\downharpoonleft\tilde{M}}+y_{\downharpoonleft K}+y_{\downharpoonleft N}$ into their $\tilde{M}-$, $K$- and $N$-parts and by tracing the ‘trajectory’ of such parts as we descend along the lower $p$-series, we can make the inclusion recorded above more precise. We choose to work modulo the $G$-invariant lattice $\tilde{M}$ and can thus focus on the $K$- and $N$-parts. The challenge is to control the contributions of $K$-parts, because they can give rise to non-trivial $K$- and $N$-parts at the next step so that the trajectory splits. In contrast, $N$-parts can only produce $N$-parts at the next step, because $N$ is $G$-invariant. At each stage along a trajectory we have, for any given $y\in L$,
$y.\mathfrak{a}_{G}$ $\in\tilde{M}\oplus(y.\mathfrak{a}_{G})_{\downharpoonleft K}\oplus(y.\mathfrak{a}_{G})_{\downharpoonleft N}$
$\subseteq\tilde{M}\oplus(y_{\downharpoonleft K}.\mathfrak{a}_{G})_{\downharpoonleft K}\oplus(y_{\downharpoonleft K}.\mathfrak{a}_{G})_{\downharpoonleft N}\oplus(y_{\downharpoonleft N}.\mathfrak{a}_{G}).$
We introduce a parameter $j$, which indicates that we follow the $K$-parts for $j$ steps, then switch to $N$-parts and continue to follow the latter for the remaining ij1i-j-1 steps. For instance, for $i=4$ and $j=2$, we would consider, for any given $y\in K$ and $a_{1},\ldots,a_{4}\in\mathfrak{a}_{G}$, the sequence of elements
$y_{0}=y,\quad y_{1}=(y_{0}.a_{1})_{\downharpoonleft K},\quad y_{2}=(y_{1}.a_{2})_{\downharpoonleft K},\quad y_{3}=(y_{2}.a_{3})_{\downharpoonleft N},\quad y_{4}=y_{3}.a_{4}.$
Concretely, we deduce that for all $i\in\mathbb{N}_{0}$,
$K.\mathfrak{a}_{G}^{\,i}$ $\subseteq\tilde{M}\oplus(K.\mathfrak{a}_{G}^{\,i})_{\downharpoonleft K}\oplus\sum_{j=0}^{i-1}\bigl{(}\bigl{(}\bigl{(}(K.\mathfrak{a}_{G}^{\,j})_{\downharpoonleft K}\bigr{)}.\mathfrak{a}_{G}\bigr{)}_{\!\downharpoonleft N}\bigr{)}.\mathfrak{a}_{G}^{\,i-1-j}$
$\subseteq\tilde{M}\oplus p^{\lfloor i\vartheta\rfloor-b_{1}}K\oplus\sum_{j=0}^{i-1}p^{\lfloor j\vartheta\rfloor-b_{1}}N.\mathfrak{a}_{G}^{\,i-1-j}$
$\subseteq\tilde{M}\oplus p^{\lfloor i\vartheta\rfloor-b_{1}}K\oplus\sum_{j=0}^{i-1}p^{\lfloor j\vartheta\rfloor+\lfloor(i-1-j)\eta\rfloor-2b_{1}}N$
$\subseteq\tilde{M}\oplus p^{\lfloor i\vartheta\rfloor-b_{1}}K\oplus p^{\min\{\lfloor(i-1)\vartheta\rfloor,\lfloor(i-1)\eta\rfloor\}-2b_{1}-2}N.$
This yields, for $i\in\mathbb{N}_{0}$,
$L_{i}=\tilde{M}.\mathfrak{a}_{G}^{\,i}+K.\mathfrak{a}_{G}^{\,i}+N.\mathfrak{a}_{G}^{\,i}\subseteq M+K.\mathfrak{a}_{G}^{\,i}+N.\mathfrak{a}_{G}^{\,i}\\ \subseteq M+K.\mathfrak{a}_{G}^{\,i}+p^{\lfloor i\eta\rfloor-b_{1}}N\subseteq M\oplus p^{\min\{\lfloor i\vartheta\rfloor,\lfloor i\eta\rfloor\}-2b_{1}-3}N.$
Thus (2.5) holds for $b_{0}=2b_{1}+3$.
Next we establish that there exist
$b,c\in\mathbb{N},\quad n_{i}\in\mathbb{N}_{0}\text{ with }\min\{\lfloor i\vartheta\rfloor,\lfloor i\eta\rfloor\}-b\leq n_{i}\leq\lfloor i\eta\rfloor\\ \text{and}\quad w_{i,k}\in p^{\min\{\lfloor i\vartheta\rfloor,\lfloor i\eta\rfloor\}-b}N,\quad\text{for $i\in\mathbb{N}_{0}$ and $\widetilde{e}<k\leq e$,}$
such that the lower $p$-series $L_{i}$, $i\in\mathbb{N}_{0}$, is $c$-equivalent to the filtration series
$\bigoplus_{k=1}^{\widetilde{e}}p^{\lfloor i\xi_{k}\rfloor}\mathbb{Z}_{p}x_{k}\;\oplus\;\bigoplus_{k=\widetilde{e}+1}^{e}\mathbb{Z}_{p}\big{(}p^{\lfloor i\xi_{k}\rfloor}x_{k}+w_{i,k}\big{)}\;\oplus\;p^{n_{i}}N,\quad i\in\mathbb{N}_{0}.$ (2.7)
@@
Moreover, the $G$-invariant $\mathbb{Z}_{p}$-lattice $\tilde{M}=\bigoplus_{k=1}^{\widetilde{e}}\mathbb{Z}_{p}x_{k}$ embeds isomorphically as $(\tilde{M}+N)/N$ into $L/N$. By Proposition 2.4, applied to $L/N$, there exists $c_{2}\in\mathbb{N}$ such that for all $i\in\mathbb{N}_{0}$,
$p^{c_{2}}\bigoplus_{k=1}^{\widetilde{e}}p^{\lfloor i\xi_{k}\rfloor}\mathbb{Z}_{p}x_{k}\subseteq\lambda_{i}(\tilde{M})\subseteq L_{i}.$
@@
$p^{c_{3}}p^{\lfloor i\xi_{k}\rfloor}x_{k}\in L_{i}\qquad\text{for $1\leq k\leq\widetilde{e}$.}$
@@
allows us to select $w_{i,\widetilde{e}+1},\ldots,w_{i,e}\in N$ such that
$p^{c_{3}}\big{(}p^{\lfloor i\xi_{k}\rfloor}x_{k}+w_{i,k}\big{)}\in L_{i}\qquad\text{for $\widetilde{e}<k\leq e$.}$
We set $b=b_{0}+c_{3}$ and deduce from (2.5) that
$w_{i,k}\in p^{\min\{\lfloor i\vartheta\rfloor,\lfloor i\eta\rfloor\}-b}N\qquad\text{for $i\in\mathbb{N}_{0}$ with $i\geq j$ and $\widetilde{e}<k\leq e$.}$
For $i\in\mathbb{N}_{0}$ with $i<j$ we set $w_{i,\widetilde{e}+1}=\ldots=w_{i,e}=0$; small values of $i$ play an insignificant role and this simple choice has no relevance beyond the fact that the series in (2.7) is properly defined. To streamline the notation, we also set
$w_{i,1}=\ldots=w_{i,\widetilde{e}}=0\qquad\text{for all $i\in\mathbb{N}_{0}$.}$ (2.8)
@@
Choosing $c_{4}\in\mathbb{N}$ with $c_{4}\geq c_{1}+c_{3}$ sufficiently large to cover also the finitely many terms indexed by $i<j$, we have established that $L_{i}$, $i\in\mathbb{N}_{0}$, is $c_{4}$-equivalent to the filtration series
@@
In view of (2.5) and $b\geq b_{0}$, we deduce from our choice of $\eta$, in connection with the lower $p$-series of $N$, that there is $c_{5}\in\mathbb{N}$ such that
@@
$p^{\lfloor i\eta\rfloor+c_{5}}N\subseteq\lambda_{i}(N)\subseteq L_{i}\cap N\subseteq p^{\min\{\lfloor i\vartheta\rfloor,\lfloor i\eta\rfloor\}-b}N\qquad\text{for $i\in\mathbb{N}_{0}$.}$
Hence Proposition 2.2 shows that, for suitable $c_{6}\in\mathbb{N}$ with $c_{6}\geq c_{5}$, the filtration series $L_{i}\cap N$, $i\in\mathbb{N}_{0}$, of $N$ is $c_{6}$-equivalent to a filtration series of the form $p^{n_{i}}N$, $i\in\mathbb{N}_{0}$, where
$\min\{\lfloor i\vartheta\rfloor,\lfloor i\eta\rfloor\}-b\leq n_{i}\leq\lfloor i\eta\rfloor\qquad\text{for all $i\in\mathbb{N}_{0}$.}$
Recalling the notational ‘nullnummer’ (2.8), we deduce that, for $c=c_{4}+c_{6}$, the lower $p$-series $L_{i}$, $i\in\mathbb{N}_{0}$, is $c$-equivalent to the filtration series in (2.7).
At this stage we are ready to finish the proof in the case that $\eta\leq\vartheta$. Indeed, in this situation
@@
for all relevant indices $i,k$ and hence a small adjustment of the series in (2.7) shows that $L_{i}$, $i\in\mathbb{N}_{0}$, is $(b+c)$-equivalent to the series
$\bigoplus_{k=1}^{\widetilde{e}}p^{\lfloor i\xi_{k}\rfloor}\mathbb{Z}_{p}x_{k}\;\oplus\;\bigoplus_{k=\widetilde{e}+1}^{e}\mathbb{Z}_{p}p^{\lfloor i\xi_{k}\rfloor}x_{k}\;\oplus\;p^{\lfloor i\eta\rfloor}N,\quad i\in\mathbb{N}_{0}.$
@@
and, if $k\in\{1,\ldots,\widetilde{e}\}$ is minimal subject to $\eta\leq\xi_{k}$, this yields a stratification with frame
@@
Below we show the following: there exists $m\in\mathbb{N}_{0}$ such that for every choice of $\mathbf{v}=(v_{\widetilde{e}+1},\ldots,v_{\widehat{e}})\in\big{(}p^{-b}N\big{)}^{\widehat{e}-\widetilde{e}}$, we have
$p^{m}N\subseteq\tilde{M}+\sum_{k=\widetilde{e}+1}^{\widehat{e}}(x_{k}+v_{k}).\mathbb{Z}_{p}G.$ (2.9)
From this we deduce that for $i\in\mathbb{N}_{0}$,
$L_{i}\supseteq p^{c}\Big{(}p^{\lfloor i\vartheta\rfloor}\tilde{M}+\bigoplus_{k=\widetilde{e}+1}^{{\widehat{e}}}p^{\lfloor i\vartheta\rfloor}\mathbb{Z}_{p}(x_{k}+v_{i,k})\Big{)}.\mathbb{Z}_{p}G\\ \supseteq p^{c+\lfloor i\vartheta\rfloor}\Big{(}\tilde{M}+\sum_{k=\widetilde{e}+1}^{\widehat{e}}(x_{k}+v_{i,k}).\mathbb{Z}_{p}G\Big{)}\supseteq p^{c+m+\lfloor i\vartheta\rfloor}N,$
hence $p^{n_{i}}N\supseteq p^{c}L_{i}\cap N\supseteq p^{2c+m+\lfloor i\vartheta\rfloor}N$, and we obtain
$w_{i,k}\in p^{\lfloor i\vartheta\rfloor-b}N\quad\text{and}\quad\lfloor i\vartheta\rfloor-b\leq n_{i}\leq\lfloor i\vartheta\rfloor+2c+m$
for all relevant indices $i,k$. Adjusting the series in (2.7), we deduce that $L_{i}$, $i\in\mathbb{N}_{0}$, is $(b+3c+m)$-equivalent to the series
$\bigoplus_{k=1}^{\widetilde{e}}p^{\lfloor i\xi_{k}\rfloor}\mathbb{Z}_{p}x_{k}\;\oplus\;\bigoplus_{k=\widetilde{e}+1}^{e}p^{\lfloor i\xi_{k}\rfloor}\mathbb{Z}_{p}x_{k}\;\oplus\;p^{\lfloor i\vartheta\rfloor}N,\quad i\in\mathbb{N}_{0}.$
@@
$(x_{1},x_{2},\ldots,x_{\widetilde{e}},y_{1},y_{2},\ldots,y_{d-e},x_{\widetilde{e}+1},x_{\widetilde{e}+2},\ldots,x_{e})$
@@
$(\xi_{1},\xi_{2},\ldots,\xi_{\widetilde{e}},\underbrace{\vartheta,\vartheta,\ldots,\vartheta}_{\text{$d-e$ entries}},\xi_{\widetilde{e}+1},\xi_{\widetilde{e}+2},\ldots,\xi_{e})\in\big{(}[\nicefrac{{1}}{{\max\{e,d-e\}}},1]\cap\mathbb{Q}\big{)}^{d}.$
It remains to justify (2.9), and for conciseness we put $\mathcal{V}=\big{(}p^{-b}N\big{)}^{\widehat{e}-\widetilde{e}}$.
For each $\mathbf{v}=(v_{\widetilde{e}+1},\ldots,v_{\widehat{e}})\in\mathcal{V}$, we think of $\tilde{T}_{\mathbf{v}}=\sum_{k=\widetilde{e}+1}^{\widehat{e}}\mathbb{Z}_{p}(x_{k}+v_{k})$ as a deformation of $T$ inside $M\oplus p^{-b}N\subseteq\mathbb{Q}_{p}\otimes_{\mathbb{Z}_{p}}L$. Recall that, by our choice of $\widetilde{e}$, the $\mathbb{Z}_{p}$-lattice $\tilde{M}\oplus p^{b}\tilde{T}_{\mathbf{v}}\subseteq L$ is not $G$-invariant, whereas $\tilde{M}\oplus p^{b}\tilde{T}_{\mathbf{v}}\oplus N=\tilde{M}\oplus p^{b}T\oplus N$ is $G$-invariant by Proposition 2.4. This implies that
$\big{(}\tilde{M}+\tilde{T}_{\mathbf{v}}.\mathbb{Z}_{p}G\big{)}\cap N\neq\{0\}.$
Clearly, the function $\mathcal{V}\to\mathbb{N}_{0}$, $\mathbf{v}\mapsto l_{\mathbf{v}}$ is locally constant and, since the domain $\mathcal{V}$ is compact, the maximum $m=\max\{l_{\mathbf{v}}\mid\mathbf{v}\in\mathcal{V}\}$ exists and has the property described in (2.9). ∎
Remark 2.8.
From the proof of Theorem 1.1 one can extract an inductive procedure for determining the growth rate $(\xi_{1},\ldots,\xi_{d})$ of a stratification for the lower $p$-series of the $\mathbb{Z}_{p}G$-module $L$.
[…]
We give a small explicit example to illustrate possible pitfalls. Let $\mathfrak{o}$ be the valuation ring of the local field $\mathbb{Q}_{p}(\pi)$, where $\pi^{4}=p$, so that $\mathfrak{o}=\mathbb{Z}_{p}+\mathbb{Z}_{p}\pi+\mathbb{Z}_{p}\pi^{2}+\mathbb{Z}_{p}\pi^{3}$ is a $\mathbb{Z}_{p}$-lattice of rank $4$. We consider the $16$-dimensional $\mathbb{Z}_{p}$-lattice $L=\bigoplus_{i=1}^{4}\mathfrak{o}.x_{i}\cong\mathfrak{o}^{4}$, equipped with the natural right-action of the pro-$p$ group $G\leq\mathsf{GL}_{4}(\mathfrak{o})$ generated by the four elements
$\begin{pmatrix}\boldsymbol{1+\pi}&0&0&0\\ 0&\boldsymbol{1}&0&0\\ 0&0&\boldsymbol{1}&0\\ 0&0&0&\boldsymbol{1}\end{pmatrix},\begin{pmatrix}\boldsymbol{1}&\boldsymbol{1}&0&0\\ 0&\boldsymbol{1}&0&0\\ 0&0&\boldsymbol{1}&0\\ 0&0&0&\boldsymbol{1}\end{pmatrix},\begin{pmatrix}\boldsymbol{1}&0&0&0\\ 0&\boldsymbol{1}&0&0\\ 0&0&\boldsymbol{1+\pi^{2}}&0\\ 0&0&0&\boldsymbol{1}\end{pmatrix},\begin{pmatrix}\boldsymbol{1}&0&0&0\\ 0&\boldsymbol{1}&0&0\\ 0&0&\boldsymbol{1}&\boldsymbol{1}\\ 0&0&0&\boldsymbol{1}\end{pmatrix}$.
It is not difficult to work out the lower $p$-series of the $\mathbb{Z}_{p}G$-module $L$ and to deduce that is has a stratification of growth rate
$(\underbrace{1/4,\ldots,1/4}_{\text{8 entries}},\underbrace{1/2,\ldots,1/2}_{\text{8 entries}}).$
Moreover, the $\mathbb{Z}_{p}G$-submodules $N$ of $L$ such that $\mathbb{Q}_{p}\otimes_{\mathbb{Z}_{p}}N$ is simple are precisely the $1$-dimensional $\mathbb{Z}_{p}$-submodules of $\{y\in L\mid\forall g\in G:y^{g}=y\}=\mathfrak{o}x_{2}+\mathfrak{o}x_{4}$. For any such a module $N$ we can arrange that $N$ is a direct summand of the $\mathbb{Z}_{p}$-lattice $L$, simply by passing to the isolator of $N$ in $L$. Furthermore, $N$ has growth rate $(\eta)=(1)$, and $L/N$ has one of the following two possible growth rates
$(\underbrace{1/4,\ldots,1/4}_{\text{8 entries}},\underbrace{1/2,\ldots.1/2}_{\text{7 entries}}),\qquad(\underbrace{1/4,\ldots,1/4}_{\text{7 entries}},\underbrace{1/2,\ldots.1/2}_{\text{8 entries}}).$
Depending on how $N$ lies within $L$, the relevant value of $\vartheta$ is $1/2$ or $1/4$.
@@
Furthermore, for every $i\in\mathbb{N}$ with $i\geq j$, the powerfully embedded subgroup $P_{i}(G)\trianglelefteq_{\mathrm{o}}U$ gives, on the Lie lattice side, a powerfully embedded Lie sublattice $L_{i^{*}}\trianglelefteq_{\mathrm{o}}L$, where we write $i^{*}=i-j$ to ease notationally a recurrent constant shift by $j$ in the index; see [16, Thm. 1.4].